# SpecServe: Efficient and SLO-Aware Large Language Model Serving with Adaptive Speculative Decoding

### v1

@@ Abstract
Large Language Model (LLM) services often face challenges in achieving low inference latency and meeting Service Level Objectives (SLOs) under dynamic request patterns. Speculative decoding, which exploits lightweight models for drafting and LLMs for verification, has emerged as a compelling technique to accelerate LLM inference. However, existing speculative decoding solutions often fail to adapt to varying workloads and system environments, resulting in performance variability and SLO violations. In this paper, we introduce SpecServe, an efficient LLM inference system that dynamically adjusts speculative strategies according to real-time request loads and system configurations. SpecServe proposes a theoretical model to understand and predict the efficiency of speculative decoding across diverse scenarios. Additionally, it implements intelligent drafting and verification algorithms to guarantee optimal performance while achieving high SLO attainment. Experimental results on real-world LLM traces demonstrate that SpecServe consistently meets SLOs and achieves substantial performance improvements, yielding 1.14$\times$-14.3$\times$ speedups over state-of-the-art speculative inference systems.
1 Introduction
@@
Large Language Models (LLMs) have demonstrated their versatility across a broad spectrum of tasks [3], such as translation [24], question answering [16], and mathematical reasoning [7]. This success has led to the rapid adoption of LLM-based services, which accept users’ real-time queries, run LLM inference, and return generated responses. During LLM inference, the decoding phase generates output tokens sequentially (i.e., autoregressive decoding), which typically dominates end-to-end request latency. Hence LLM inference systems need to meet stringent performance requirements for token generation—often specified as Service-Level Objectives (SLOs) [40].
Speculative decoding has recently emerged as an appealing solution for accelerating LLM inference [18, 5]. It leverages a lightweight draft model to generate multiple draft tokens and then runs the target model (i.e., LLM) to verify them all at once. Compared with autoregressive decoding, this approach guarantees the same output results without needing to load LLMs for every token generation, effectively improving GPU utilization and reducing inference latency.
@@
However, existing speculative decoding solutions often suffer from suboptimal performance when dealing with dynamic request patterns and varying system configurations, which are typical scenarios in real-world online LLM services. These solutions generally employ a pre-defined speculative length, i.e., the number of draft tokens generated in each round. While a larger speculative length increases the chance to accept more valid tokens in verification, it incurs higher latency and risks resource wastage due to potential token rejections [22]. Additionally, we observe that the optimal configuration of speculative length depends on the interplay between request loads, workload categories, and hardware resources (see §2.3). Consequently, using a pre-defined speculative length fails to consistently deliver satisfactory performance under changing workloads and environments.
Moreover, current speculative decoding solutions often do not take into account token-level SLOs of LLM inference, failing to ensure high SLO attainment. Although speculative decoding can achieve significantly lower inference latencies compared to autoregressive decoding, the uncertain token acceptance during verifications can lead to considerable fluctuations in overall token generation latency, such as Time Per Output Token (TPOT). This variability makes it challenging for speculative decoding to consistently meet SLO requirements, thereby affecting the ability to provide users with consistent and efficient services.
@@
Therefore, we advocate that an efficient speculative inference system should enable fine-grained control over speculative lengths and beware of dynamic workloads and SLO requirements. Following this idea, we present SpecServe, a system that employs adaptive speculative decoding to reduce the end-to-end latency of online LLM inference. It dynamically determines the optimal speculative length for varying request patterns, system configurations, and SLO constraints. By accommodating these changing conditions, SpecServe strives to achieve lower inference latency and ensure SLO attainment. However, achieving adaptive speculative decoding in SpecServe presents three key challenges.
The first challenge is to build an efficiency model for speculative decoding that reflects dynamic system loads and diversified request categories. In online LLM inference, system load can be simply represented by metrics such as batch size and context length, but it is challenging to measure the impact of request complexity on the efficiency of speculative decoding. This is because it is not possible to know during the drafting phase whether the draft tokens will be accepted in the verification stage. Based on our observations, the acceptance rate of draft tokens is directly proportional to the confidence score at which they are generated by the draft model [8, 20]. Therefore, we propose the adaptive drafter, which integrates the efficiency model of speculative decoding into the execution of the draft model. It directly uses the confidence score of draft tokens to update their acceptance rate in real-time. Then, it employs a simple yet effective linear model to depict the impact of system load and system configuration on the execution time of speculative decoding. Ultimately, the optimal speculative length is determined collaboratively by the acceptance rate of draft tokens and the execution time of speculative decoding.
Secondly, SpecServe requires fine-grained control over the speculative length to deliver optimal inference performance. Due to the diversity of request categories, applying the same speculative length to different requests within the same batch still fails to achieve optimal latency. Therefore, we propose a succinct and efficient solution, named the confidence prior verifier. The essence of this solution lies in decoupling the drafting and verification phases, and leveraging the acceptance rate of draft tokens recorded by the adaptive drafter, as well as a speculative decoding execution time model. Through this approach, we can significantly reduce resource wastage caused by verifying useless draft tokens, by only allowing the target model to verify those tokens that are more likely to be accepted, without differentiating the speculative length of each request.
The third challenge is to improve SLO attainment for online LLM inference services. The SLO violations in speculative decoding originate from the discrepancy between the expected and actual number of accepted tokens, indicating that the additional decoding cost does not correspond to a proportional increase in generated tokens. To mitigate the issue of SLO violations, we leverage SLO constraints to control the speculative length. We integrate these SLO constraints directly into the efficiency model of speculative decoding, resulting in the development of a SLO-aware efficiency estimator for speculative decoding. This SLO-aware efficiency estimator considers not only the speed of token generation but also the quality of these tokens in meeting predefined SLO constraints.
@@
We have implemented SpecServe atop vLLM [17] and evaluated it using real-world LLM inference traces. Evaluation results show that SpecServe can effectively adapt to varying workloads and system configurations, consistently achieving low inference latency and high SLO attainments. Compared with state-of-the-art speculative inference systems, SpecServe enables 1.14$\times$ to 14.3$\times$ speedups while meeting SLO requirements.
@@
  • We proposed SpecServe, an efficient LLM inference system with adaptive speculative decoding.
@@
  • We evaluated SpecServe using real-world production traces, which demonstrates that SpecServe can significantly improve inference performance by up to 14.3 $\times$ compared with baselines.
2 Background and Motivation
2.1 LLM Inference Service
Figure 1: Different request patterns in real production traces.
@@ 2.1 LLM Inference Service
In LLM inference [1, 2], the model accepts a list of tokens $\mathbf{x}=(x_{1},\ldots,x_{n})$, and outputs the conditional probability distribution of the next token $(x_{n+1}|x_{1},\ldots,x_{n})$. The trained LLMs are typically deployed onto inference systems to provide online services, where they receive requests consisting of token lists and generate a complete output sequence $(x_{n+1},\ldots,x_{n+T})$ by predicting the next token $T$ times.
@@
In real-world LLM inference systems, they usually need to handle dynamic request arrivals from a variety of workloads. Fig. 1 illustrates two representative LLM inference traces from Azure [29]. The first pattern (trace 1) demonstrates sporadic bursts of requests, characterized by a generally low request rate but with extremely high peaks; the second pattern (trace 2) displays lower request volatility, yet still reflects the dynamic nature of system load. A key requirement of LLM inference is to meet stringent latency requirements (i.e., SLOs), which necessitates inference systems to efficiently handle dynamic requests and achieve low-latency token generation [10, 40, 41, 39].
@@
Request batching is a common approach to improve inference performance and resource efficiency for LLM inference [38, 12]. It fuses multiple requests associated with the same model as a batch, and executes them together to fully exploit parallel computing resources of hardware accelerators such as GPUs. Mainstream LLM inference systems such as TensorRT-LLM [28] and vLLM [17] offer the ability of batching for enhanced efficiency. While request batching enhances GPU utilization during decoding, this method fails to address inefficiencies in scenarios with persistently low per-model request rates. According to real-world inference traces (Fig. 2), LLM services frequently experience prolonged periods of low request rates—a pattern consistent with prior observations of inference workloads[39]. These findings underscore a critical requirement for LLM inference systems: maintaining high computational efficiency even under sustained low request arrival rates.
2.2 Speculative Decoding
@@
Figure 2: A single-step speculative decoding process with a speculative length of 1. We primarily discuss the greedy sampling, where the token with the highest confidence score is selected as the output instead of the whole distribution.
@@
Speculative decoding addresses the inefficiency of traditional autoregressive decoding in LLM inference systems. This approach utilizes a small draft model to predict the following tokens and verify them at once, improving resource efficiency and end-to-end performance. As shown in Fig. 2, speculative decoding can be divided into two parts: drafting and verification. During the drafting phase, given a list of tokens $\mathbf{x}=(x_{1},\ldots,x_{n})$, speculative decoding uses an efficient draft model to generate a draft sequence $\mathbf{y}=(y_{1},\ldots,y_{SL})$, where $SL$ is the speculative length. In the verification process, the draft sequence $\mathbf{y}=(y_{1},\ldots,y_{SL})$ is input into the target model in parallel for a single forward pass, generating multiple output distributions $P(y_{1}|x_{1:n}),\ldots,P(y_{SL}|x_{1:n},y_{1:SL-1})$. Based on these output distributions and specific acceptance criteria, all tokens $(y_{1},\ldots,y_{m-1})$ before the rejected token $y_{m}$ that first fails the acceptance criteria will be selected as accepted tokens, plus a bonus token $z_{m}$ resampled from distribution $P(y_{m}|x_{1:n},y_{1:m-1})$, to form the output sequence $(y_{1},\ldots,y_{m-1},z_{m})$ generated by speculative decoding. Since the acceptance of tokens is still determined by the output distribution of the target model, speculative decoding is lossless compared to auto-regressive decoding.
@@
The speculative length, by controlling the number of candidate tokens generated by the draft model, directly impacts both the latency benefits and resource consumption of speculative decoding. Current mainstream methods [18, 33, 26, 4, 19] rely on offline analysis to determine a static speculative length, with its optimal value typically identified through heuristic search on specific datasets and under fixed hardware configurations. Furthermore, while a few methods, such as SmartSpec [22], are capable of setting the speculative length for each decoding step, their coarse-grained control still fails to adapt to the fine-grained variations present in real-world request scenarios.
2.3 Limitations of Existing Solutions
@@
Current speculative decoding approaches face three fundamental limitations in the practical deployment of LLM inference systems. These limitations stem from their lack of adaptability to dynamic request patterns, system configurations, and SLO constraints.
@@
Figure 3: Speedup across various speculative lengths under diverse request quantities and types. The red dashed line represents the speedup value of 1.
@@
Limitation #1: Inefficiency under dynamic workloads. As shown in Fig. 3, conventional speculative decoding methods, which rely on fixed speculative lengths, fail to adapt to dynamic changes in request patterns. Under high-load scenarios (batch size 128), excessively long speculative lengths lead to: (1) increased computational overhead during the verification phase due to redundant tokens generated by the draft model; (2) resource contention caused by batch sizes exceeding device memory. Conversely, under low-load scenarios (batch size 1), excessively short speculative lengths fail to fully utilize idle computational resources, leaving the acceleration potential of speculative decoding untapped. Workload characteristics represent another critical variable. We demonstrate this by analyzing workloads extracted from RAG [14], mathematical reasoning [7], and question-answering [16] datasets, each exhibiting different optimal speculative lengths. Evidently, applying the same speculative length across different requests results in both redundant token generation and wasted acceleration potential, making it difficult to achieve optimal latency optimization.
@@
Figure 4: Speedup across different speculative lengths for various draft-target model pairs on diverse computing platforms.
@@
Limitation #2: Agnostic to system configurations. As illustrated in Fig. 4, when hardware configurations or model pairings change, the fixed-length strategy of traditional speculative decoding fails to maintain optimality. This phenomenon arises from changes in the hardware roofline model: when the arithmetic intensity of the target model is below the platform’s balance point, parallel token processing during the verification phase can effectively improve computational efficiency [25]. Conversely, in compute-bound scenarios, the probabilistic nature of speculative decoding reduces its token generation efficiency compared to autoregressive decoding, resulting in latency degradation.
@@
Limitation #3: SLO violations. As revealed in Fig. 5, while speculative decoding can achieve higher acceleration ratios compared to autoregressive decoding, its SLO attainment may be significantly lower. This issue arises from the probabilistic variability of speculative decoding’s acceleration effects: longer speculative lengths increase the expected acceleration ratio but also raise the risk of exceeding the TPOT limit (e.g., when the number of tokens actually generated is lower than expected, the TPOT for that speculative decoding attempt will significantly exceed its expected value). This uncertainty prevents speculative decoding from simultaneously meeting the high-performance and high-reliability requirements of inference systems.
2.4 Key Insight and Challenges
Figure 5: Relationship between SLO attainment and end-to-end latency speedup under different speculative lengths. The red dashed line represents the speedup value of 1 and the SLO attainment value of 0.9.

### v2

# AdaSpec: Adaptive Speculative Decoding for Fast, SLO-Aware Large Language Model Serving

@@ Abstract.
Cloud-based Large Language Model (LLM) services often face challenges in achieving low inference latency and meeting Service Level Objectives (SLOs) under dynamic request patterns. Speculative decoding, which exploits lightweight models for drafting and LLMs for verification, has emerged as a compelling technique to accelerate LLM inference. However, existing speculative decoding solutions often fail to adapt to fluctuating workloads and dynamic system environments, resulting in impaired performance and SLO violations. In this paper, we introduce AdaSpec, an efficient LLM inference system that dynamically adjusts speculative strategies according to real-time request loads and system configurations. AdaSpec proposes a theoretical model to analyze and predict the efficiency of speculative strategies across diverse scenarios. Additionally, it implements intelligent drafting and verification algorithms to maximize performance while ensuring high SLO attainment. Experimental results on real-world LLM service traces demonstrate that AdaSpec consistently meets SLOs and achieves substantial performance improvements, delivering up to 66% speedup compared to state-of-the-art speculative inference systems. The source code is publicly available at https://github.com/cerebellumking/AdaSpec
large language models, machine learning inference, speculative decoding
*Corresponding authors. E-mail addresses: yuminchen@cuhk.edu.cn (Minchen Yu), shiqj@tongji.edu.cn (Qingjiang Shi)..
journalyear: 2025copyright: acmlicensedconference: ACM Symposium on Cloud Computing; November 19–21, 2025; Online, USAbooktitle: ACM Symposium on Cloud Computing (SoCC ’25), November 19–21, 2025, Online, USAdoi: 10.1145/3772052.3772239isbn: 979-8-4007-2276-9/25/11submissionid: 179ccs: Computer systems organization Cloud computingccs: Computer systems organization Parallel architectures
1. Introduction
@@
Large Language Models (LLMs) have demonstrated remarkable versatility across a broad spectrum of tasks (Brown et al., 2020), such as translation (Lu et al., 2024), question answering (Kwiatkowski et al., 2019), and mathematical reasoning (Cobbe et al., 2021). This success has fueled the rapid and widespread adoption of cloud-based LLM services, which enable users to submit real-time queries and receive generated responses via LLM inference. The inference process of LLMs typically consists of two main phases: the prefill phase, where the model processes the user’s input prompt and initializes the context, and the decoding phase, where the LLM generates output tokens sequentially. The decoding phase generally dominates the end-to-end request latency, particularly for long output sequences. Therefore, a critical requirement for LLM services is to guarantee fast and consistent token generation performance, which is often specified as Service-Level Objectives (SLOs) such as a maximum allowable Time Per Output Token (TPOT) (Patke et al., 2024; Hu et al., 2021).
Speculative decoding has recently emerged as a compelling solution for accelerating LLM inference (Leviathan et al., 2023; Chen et al., 2023). This approach leverages a lightweight draft model to produce multiple candidate tokens, which are then collectively verified by the target LLM. By generating and validating multiple tokens at once, speculative decoding produces the same outputs as traditional autoregressive decoding, but greatly reduces the overhead associated with invoking the LLM for each token. This efficiency leads to improved GPU utilization, lower inference latency, and higher overall system throughput.
@@
However, existing speculative decoding solutions often face notable challenges in real-world cloud scenarios, where request patterns and resource configurations are highly dynamic. Most current solutions rely on a pre-defined speculative length—the number of draft tokens generated per speculative round. While increasing this speculative length can raise the likelihood of verifying more tokens in a single pass, it also amplifies the risk of token rejections, leading to wasted computation and higher end-to-end latency (Liu et al., 2024b). Finding the optimal speculative length is complicated, as it depends on a complex interplay of factors such as request rates, model loads, and hardware resources (see §2.3). Consequently, static approaches that use a fixed speculative length fails to consistently deliver satisfactory performance under changing workloads and environments.
Moreover, existing solutions often overlook token-level SLOs in LLM inference, leading to inconsistent SLO attainment. Although speculative decoding reduces inference latencies compared to the autoregressive approach, the variability in token acceptance during the verification process can cause significant fluctuations in overall token generation latency, such as TPOT. This inherent variability poses challenges for speculative decoding to consistently satisfy SLO requirements, undermining the ability to deliver reliable and predictable quality of service to users.
@@
To address these limitations, an efficient speculative inference system should enable fine-grained, adaptive control over speculative lengths, remaining responsive to dynamic workloads and SLO requirements. Following this principle, we present AdaSpec, a system that leverages adaptive speculative decoding to improve performance and SLO attainment for LLM inference. AdaSpec aims to dynamically configure the optimal speculative length for each request, adapting to fluctuating request patterns, diverse system configurations, and varying SLO constraints. However, achieving adaptive speculative decoding in AdaSpec presents three challenges.
The first challenge is to accurately characterize the efficiency of speculative decoding under real-world, dynamic conditions. In cloud environments, LLM inference workloads can vary significantly in batch size, context length, and workload heterogeneity. While system-level characteristics such as batch size and context length are easy to monitor, the acceptance rate of draft tokens is difficult to assess during drafting. This challenge stems from the inherent uncertainty: we can only detect whether drafted tokens will be accepted until the verification phase. Without a reliable efficiency estimation mechanism, selecting the optimal speculative length becomes infeasible. To address this, AdaSpec introduces an efficiency model (see §4) that estimates both the expected acceptance rate and execution time, leveraging real-time draft token confidence scores and current system configurations. This model empowers the adaptive drafter (see §5.1) to dynamically adjust the speculative length at the batch level. Through this approach, AdaSpec achieves more informed and responsive control over speculative decoding, significantly enhancing overall system performance.
Second, enabling fine-grained and flexible control of speculative length at the request level presents significant challenges. In practice, LLM inference systems batch heterogeneous user requests to maximize throughput and resource utilization. However, existing speculative decoding methods generally apply a uniform speculative length for all requests within a batch, overlooking substantial differences in token acceptance probabilities among queries. This rigid approach can result in inefficiencies—for example, applying the same speculative length for both simple and complex prompts may cause unnecessary token rejections or insufficient acceleration for less demanding requests. To tackle this challenge, AdaSpec introduces the confidence prior verifier (see §5.2), which decouples the drafting and verification stages of decoding. By leveraging per-request acceptance probability estimates, the system can selectively allocate verification efforts to tokens with higher predicted acceptance rates. This enables each request in a batch to effectively utilize its own optimal speculative length, without requiring changes to the overall batch execution pipeline.
Third, AdaSpec must balance the efficiency gains of speculative decoding with the need to meet SLOs. Speculative decoding introduces probabilistic performance variations: while longer speculative lengths can boost throughput, they also increase the likelihood of token rejections and latency spikes, which may result in SLO violations. Existing approaches often neglect this crucial trade-off, leading to degraded service quality under strict latency constraints. AdaSpec addresses this by incorporating a SLO-aware efficiency estimator (see §5.3) that proactively evaluates whether a chosen speculative decoding configuration can satisfy TPOT requirements before execution. This enables AdaSpec to make adaptive, real-time decisions that optimize for both high performance and SLO compliance, ensuring robust efficiency without sacrificing reliability.
@@
We have implemented AdaSpec atop vLLM (Kwon et al., 2023) and evaluated using real-world LLM inference traces. Evaluation shows that AdaSpec can adapt to varying workloads and system configurations, consistently achieving low latency and high SLO attainments. Compared with state-of-the-art speculative inference systems, AdaSpec enables up to 66% speedups while meeting SLO requirements.
@@
  • We proposed AdaSpec, an efficient LLM inference system with adaptive speculative decoding.
@@
  • We evaluated AdaSpec using real-world production traces, which demonstrates that AdaSpec can improve inference performance by up to 66% compared with baselines.
2. Background and Motivation
2.1. LLM Inference Service
Figure 1. Different request patterns in real production traces.
@@ 2.1. LLM Inference Service
In LLM inference (Achiam et al., 2023; Bengio et al., 2000), the model accepts a list of tokens $\mathbf{x}=(x_{1},\ldots,x_{n})$, and outputs the conditional probability distribution of the next token $(x_{n+1}|x_{1},\ldots,x_{n})$. Trained LLMs are deployed onto inference systems to provide online services, where they receive requests consisting of token lists and generate a complete output sequence $(x_{n+1},\ldots,x_{n+T})$ by predicting the next token $T$ times.
@@
In real-world LLM inference systems, they usually need to handle dynamic request arrivals from a variety of workloads. Fig. 1 displays three representative LLM inference traces from BurstGPT (Wang et al., 2024) and Mooncake (Qin et al., 2025). The first pattern (trace 1) exhibits substantial request rate variations, demonstrating significant minute-level fluctuations in system load. The second pattern (trace 2) reveals high-frequency periodic oscillations, indicating batch-style request arrivals that form intensive yet regular load patterns. The third pattern (trace 3) shows rapid fluctuations around a central value, with second-level dynamics that reflect both the stochastic nature and instantaneous variability of workload. Collectively, these distinct patterns capture the diversity and complex dynamics of real-world production system loads. A key requirement of LLM inference is to meet stringent latency requirements (i.e., SLOs), which necessitates inference systems to efficiently handle dynamic requests and achieve low-latency token generation (Fu et al., 2024b; Zhang et al., 2019, 2023; Yu et al., 2024).
@@
Request batching is a common approach to improve inference performance and resource efficiency for LLM inference (Yu et al., 2022; Gao et al., 2018). It fuses multiple requests associated with the same model as a batch, and executes them together to fully exploit parallel computing resources of hardware accelerators such as GPUs. Mainstream LLM inference systems such as TensorRT-LLM (NVIDIA, 2024) and vLLM (Kwon et al., 2023) offer the ability of batching for enhanced efficiency. While request batching enhances GPU utilization during decoding, this method fails to address inefficiencies in scenarios with persistently low per-model request rates. According to real-world inference traces (Fig. 2), LLM services frequently experience prolonged periods of low request rates—a pattern consistent with prior observations of inference workloads(Yu et al., 2024). These findings underscore a critical requirement for LLM inference systems: maintaining high computational efficiency even under sustained low request arrival rates.
2.2. Speculative Decoding
@@
Figure 2. A single-step speculative decoding process with a speculative length of 1. We primarily discuss the greedy sampling, where the token with the highest confidence score is selected as the output instead of the whole distribution.
@@
Speculative decoding addresses the inefficiency of autoregressive decoding in LLM inference systems. This approach utilizes a small draft model to predict the following tokens and verify them at once, improving resource efficiency and end-to-end performance. As shown in Fig. 2, speculative decoding has two parts: drafting and verification. During the drafting phase, given a list of tokens $\mathbf{x}=(x_{1},\ldots,x_{n})$, speculative decoding uses an efficient draft model to generate a draft sequence $\mathbf{y}=(y_{1},\ldots,y_{SL})$, where $SL$ is the speculative length. In the verification process, the draft sequence is input into the target model in parallel for a single forward pass, generating multiple output distributions $P(z_{1}|x_{1:n}),\ldots,P(z_{SL+1}|x_{1:n},y_{1:SL})$. Each drafted token $y_{t}$ is compared with $z_{t}$: if $y_{t}=z_{t}$, the token is accepted; otherwise, verification stops at position $m=t$, where $z_{m}\neq y_{m}$ is the first rejected token. The final output consists of the accepted draft tokens $(y_{1},\ldots,y_{m-1})$ and one bonus token $z_{m}$ directly from the target model, forming the output $(y_{1},\ldots,y_{m-1},z_{m})$. Since all accepted tokens and bonus tokens come from the target model’s predictions, speculative decoding guarantees lossless equivalence to standard autoregressive decoding.
@@
The speculative length, by controlling the number of candidate tokens generated by the draft model, directly impacts both the latency benefits and resource consumption of speculative decoding. Mainstream methods (Leviathan et al., 2023; Sun et al., 2023; Miao et al., 2024; Cai et al., 2024; Li et al., [n. d.]) rely on offline analysis to determine a static speculative length, with its optimal value typically identified through heuristic search on specific datasets and under fixed hardware configurations. Furthermore, while a few methods, such as SmartSpec (Liu et al., 2024b), are capable of setting the speculative length for each decoding step, their coarse-grained control still fails to adapt to the fine-grained variations in real-world request scenarios.
2.3. Limitations of Existing Solutions
@@
Current speculative decoding approaches face three limitations in the practical deployment of LLM inference systems in the cloud. These limitations stem from their lack of adaptability to dynamic request patterns, system configurations, and SLO constraints.
@@
Figure 3. Speedup across various speculative lengths under diverse request quantities and types. The red dashed line represents the speedup value of 1.
@@
Limitation #1: Inefficiency under dynamic workloads. We tested the speedup ratio of speculative decoding compared to autoregressive decoding under different request quantities (batch sizes of 1, 64, 128) and different request types (retrieval-augmented generation, mathematical reasoning, question answering), with identical system configurations across various speculative lengths. As shown in Fig. 3, conventional speculative decoding methods, which rely on fixed speculative lengths, fail to adapt to dynamic changes in request patterns. Under high-load scenarios (batch size 128), excessively long speculative lengths lead to: (1) increased computational overhead during the verification phase due to redundant tokens generated by the draft model; (2) resource contention caused by batch sizes exceeding device memory. Conversely, under low-load scenarios (batch size 1), excessively short speculative lengths fail to fully utilize idle computational resources, leaving the acceleration potential of speculative decoding untapped. Workload characteristics represent another critical variable. We demonstrate this by analyzing workloads extracted from RAG (Karpukhin et al., 2020), mathematical reasoning (Cobbe et al., 2021), and question-answering (Kwiatkowski et al., 2019) datasets, each exhibiting different optimal speculative lengths. Evidently, applying the same speculative length across different requests results in both redundant token generation and wasted acceleration potential, making it difficult to achieve optimal latency optimization.
@@
Figure 4. Speedup across different speculative lengths for various draft-target model pairs on diverse computing platforms.
@@
Limitation #2: Agnostic to system configurations. We evaluated the speedup ratio of speculative decoding compared to autoregressive decoding across different draft-target model pairs deployed on varying hardware configurations (see Tab. 1) under multiple speculative lengths. As shown in Fig. 4, when hardware configurations (e.g., GPU models and counts in the computing platform) or model pairings (combinations of draft and target models) vary, the fixed-length strategy employed in traditional speculative decoding loses its optimality. The root cause lies in shifts in the hardware roofline model: when the target model’s arithmetic intensity falls below the platform’s balance point, the parallel token processing in the verification phase can significantly improve computational efficiency (Marzollo et al., 2024). However, under compute-bound conditions, the probabilistic nature of speculative decoding (where proposed draft tokens may be rejected) reduces its token generation efficiency compared to autoregressive decoding, leading to latency degradation. This configuration sensitivity highlights the method’s inability to dynamically adapt to changing system environments.
@@
Limitation #3: SLO violations. We evaluated the SLO attainment of speculative decoding under different speculative lengths, using the P90 TPOT of autoregressive decoding as the threshold. As revealed in Fig. 5, while speculative decoding can achieve higher acceleration ratios compared to autoregressive decoding, its SLO attainment may be significantly lower. This issue arises from the probabilistic variability of speculative decoding’s acceleration effects: longer speculative lengths increase the expected acceleration ratio but also raise the risk of exceeding the TPOT limit (e.g., when the number of tokens actually generated is lower than expected, the TPOT for that speculative decoding attempt will significantly exceed its expected value). This uncertainty prevents speculative decoding from simultaneously meeting the high-performance and high-reliability requirements of inference systems.
2.4. Key Insight and Challenges